OpenStudy (anonymous):

Find the circumference of the circle whose diameter is 5.95 yards

OpenStudy (anonymous):

Well, the circumference is equal to \[2 \times \pi \times r\] and D = 2r = 5.95 yards. So your circumference is just 5.95 times Pi.